# Budget-Constrained Tool Learning with Planning

### v1

@@ Abstract
Despite intensive efforts devoted to tool learning, the problem of budget-constrained tool learning, which focuses on resolving user queries within a specific budget constraint, has been widely overlooked. This paper proposes a novel method for budget-constrained tool learning. Our approach involves creating a preferable plan under the budget constraint before utilizing the tools. This plan outlines the feasible tools and the maximum number of times they can be employed, offering a comprehensive overview of the tool learning process for large language models. This allows them to allocate the budget from a broader perspective. To devise the plan without incurring significant extra costs, we suggest initially estimating the usefulness of the candidate tools based on past experience. Subsequently, we employ dynamic programming to formulate the plan. Experimental results demonstrate that our method can be integrated with various tool learning methods, significantly enhancing their effectiveness under strict budget constraints. [^1] Code can be found at https://github.com/ZYH111/BTP.
@@
[…]
@@ 1 Introduction
Despite their effectiveness, existing methods often overlook a critical aspect: utilizing tools incurs expenses, such as money and time, and users may have (implicit) budget constraints in real-world scenarios. Without meticulous management, the costs can rapidly exceed the acceptable threshold of a user. In scenarios where the expenses overshadow the benefits, even if the tool successfully addresses the intended problem, users may still perceive the solution as unsatisfactory due to the disproportionate costs incurred. Table 1 presents preliminary results for three strong tool learning approaches, ReAct (Yao et al., 2023b), DFSDT (Qin et al., 2023b), and ToT-DFS (Yao et al., 2023a), evaluated on the ToolBench (Qin et al., 2023b) dataset, with success defined strictly in terms of problem resolution within budget constraints. It is evident that all three methods experience significant declines in performance as the budget constraint tightens from unlimited to 20. Consequently, we argue that more efforts are worth devoting to the problem of budget-constrained tool learning, which aims at resolving user queries within a given budget constraint.
@@ 2.2 Tool Usage Planning
We also notice that repeatedly using the same tool too many times may be suboptimal within the budget constraint, since we expect that more diverse information can be obtained if different tools are used. Thus, we set an estimated frequency constraint $\tilde{F}(q_{u},t_{i})$ for each candidate tool $t_{i}$ and expect that $t_{i}$ should be used no more than $\tilde{F}(q_{u},t_{i})$ times. Specifically, if a tool frequently returns useless or erroneous messages, we should prevent the LLM from using it to reduce the cost of tool learning. Thus, we set a threshold $\tau$ and set $\tilde{F}(q_{u},t_{i})=0$ if $v(q_{u},t_{i})<\tau$. Otherwise, for each query $q$ in the past experience where $t_{i}$ is used in the process of resolving $q$, we count the number of times $F(q,t_{i})$ for which $t_{i}$ is used, and then calculate the weighted average of $F(q,t_{i})$ as the estimated frequency constraint $\tilde{F}(q_{u},t_{i})$:
@@ 2.4 Blacklist Mechanism
To further reduce the cost of tool learning, we introduce a blacklist mechanism to reduce the cost of repeatedly using unhelpful tools. We build the blacklist according to the judgment about the returned results. Specifically, after the LLM uses a tool and receives the returned result, we use the classifier introduced in Section 2.2 to judge whether the returned result is helpful in resolving the user query. [^4] The cost of the classifier can also be added to $c_{i}$. If a tool returns unhelpful results, we list it into the blacklist and forbid the LLM from using it. Specifically, we inform the LLM that the blacklisted tools are forbidden to use by giving the blacklisted tools in the input context during the tool learning process. If the LLM occasionally uses a blacklisted tool, we prevent the LLM from the actual action of the tool usage and return an error message.
@@ 3.2 Main Results
Method PRnormal-$\uparrow$ RFBCnormal-$\downarrow$
@@
3.3 Planned v.s. Actual Tool Usage
3.4 Effect of Budget Constraint $R$
@@ 4 Related Work
This work is also highly related to Kim et al. (2023), which use different tools in parallel to improve the efficiency of tool learning. However, their work mainly focuses on time efficiency rather than budget constraint.

### v2

@@ Abstract
Despite intensive efforts devoted to tool learning, the problem of budget-constrained tool learning, which focuses on resolving user queries within a specific budget constraint, has been widely overlooked. This paper proposes a novel method for budget-constrained tool learning. Our approach involves creating a preferable plan under the budget constraint before utilizing the tools. This plan outlines the feasible tools and the maximum number of times they can be employed, offering a comprehensive overview of the tool learning process for large language models. This allows them to allocate the budget from a broader perspective. To devise the plan without incurring significant extra costs, we suggest initially estimating the usefulness of the candidate tools based on past experience. Subsequently, we employ dynamic programming to formulate the plan. Experimental results demonstrate that our method can be integrated with various tool learning methods, significantly enhancing their effectiveness under strict budget constraints. [^1] Code can be found at https://github.com/THUNLP-MT/BTP.
{CJK}
UTF8gbsn
@@
Yuanhang Zheng1, Peng Li2∗, Ming Yan3, Ji Zhang3, Fei Huang3 and Yang Liu1,2,4thanks:   Corresponding authors: P.Li (lipeng@air.tsinghua. edu.cn) and Y.Liu (liuyang2011@tsinghua.edu.cn). 1Dept. of Comp. Sci. & Tech., Institute for AI, Tsinghua University, Beijing, China 2Institute for AI Industry Research (AIR), Tsinghua University, Beijing, China 3Institute of Intelligent Computing, Alibaba Group 4Jiangsu Collaborative Innovation Center for Language Competence, Jiangsu, China
@@ 1 Introduction
Despite their effectiveness, existing methods often overlook a critical aspect: utilizing tools incurs expenses, such as money and time, and users may have (implicit) budget constraints in real-world scenarios. Without meticulous management, the costs can rapidly exceed the acceptable threshold of a user. In scenarios where the expenses overshadow the benefits, even if the tool successfully addresses the intended problem, users may still perceive the solution as unsatisfactory due to the disproportionate incurred costs. Table 1 presents preliminary results for three strong tool learning approaches, ReAct (Yao et al., 2023b), DFSDT (Qin et al., 2023b), and ToT-DFS (Yao et al., 2023a), evaluated on the ToolBench (Qin et al., 2023b) dataset, with success defined strictly in terms of problem resolution within budget constraints. It is evident that all three methods experience significant declines in performance as the budget constraint tightens from unlimited to 20. Consequently, we argue that more efforts are worth devoting to the problem of budget-constrained tool learning, which aims at resolving user queries within a given budget constraint.
@@ 2.2 Tool Usage Planning
We also notice that repeatedly using the same tool too many times may be suboptimal within the budget constraint, since we expect that more diverse information can be obtained if different tools are used. Thus, we set an estimated frequency constraint $\tilde{F}(q_{u},t_{i})$ for each candidate tool $t_{i}$ and expect that $t_{i}$ should be used no more than $\tilde{F}(q_{u},t_{i})$ times. Specifically, if a tool frequently returns useless or erroneous messages, we should prevent the LLM from using it to reduce the cost of tool learning. Thus, we set a threshold $\tau$ and set $\tilde{F}(q_{u},t_{i})=0$ if $v(q_{u},t_{i})<\tau$ to filter out the useless tools based on the past experience. Otherwise, for each query $q$ in the past experience where $t_{i}$ is used in the process of resolving $q$, we count the number of times $F(q,t_{i})$ for which $t_{i}$ is used, and then calculate the weighted average of $F(q,t_{i})$ as the estimated frequency constraint $\tilde{F}(q_{u},t_{i})$:
@@ 2.4 Blacklist Mechanism
To further reduce the cost of tool learning, we introduce a blacklist mechanism to reduce the cost of repeatedly using unhelpful tools. During the tool learning process, we temporarily build the blacklist according to the judgment about the returned results. Specifically, after the LLM uses a tool and receives the returned result, we use the classifier introduced in Section 2.2 to judge whether the returned result is helpful in resolving the user query. [^4] The cost of the classifier can also be added to $c_{i}$. If a tool returns unhelpful results, we list it into the blacklist and forbid the LLM from using it. Specifically, we inform the LLM that the blacklisted tools are forbidden to use by giving the blacklisted tools in the input context during the tool learning process. If the LLM occasionally uses a blacklisted tool, we prevent the LLM from the actual action of the tool usage and return an error message.
@@ 3.2 Main Results
3.3 Planned v.s. Actual Tool Usage
3.4 Effect of Budget Constraint $R$
[…]
@@ 4 Related Work
This work is also highly related to Kim et al. (2023), who use different tools in parallel to improve the efficiency of tool learning. However, their work mainly focuses on time efficiency rather than budget constraint.
@@ 5 Conclusion
Acknowledgments
This work is supported by the National Natural Science Foundation of China (No. 61925601, 62276152). We thank Zonghan Yang for providing helpful suggestions about the idea of this work. We also appreciate the anonymous reviewers for their valuable comments.
@@ A.2 Further Implementation Details
Moreover, the system prompt used during in the tool learning process is constructed based on the prompt for solution path annotation introduced in Qin et al. (2023b). We add extra content presented in Figure 4 in the system prompt if some of the tools are forbidden to use. Specifically, we append the prompt Part I to the system prompt if at least one candidate tool is removed from $\mathcal{T}$. We append the prompt Part II to the system prompt if at least one candidate tool is blacklisted. We append the prompt Part III to the system prompt if all the candidate tools are forbidden to use (i.e. removed from $\mathcal{T}$ or blacklisted).
@@ A.5 Licenses of Tools, Models and Datasets
Figure 4: The extra content which may be appended to the system prompt during the tool learning process.